# Contracting with a Mechanism Designer

### v1

@@ 2.2 Model Discussion
In this subsection, we discuss key assumptions regarding modeling choice.
To justify the assumption that the principal initiates the contract, we note that in real-world crowdsourcing markets, requesters typically hold greater bargaining power because multiple intermediaries compete for their business. For example, both Upwork and Freelancer charge clients a similar fee of around 3 (Mert Bulut, 2024). This competitive dynamic is further illustrated by shifts in the microtask market: on Amazon Mechanical Turk, requesters historically benefited from low fees, but after Amazon raised its commission to 20% in 2015, many migrated to alternatives such as Prolific, which offered better user terms and higher data quality (Banakou and Slater, 2023). These observations align with standard contract theory, where principals with strong outside options propose contracts that agents accept.
@@
To justify the assumption in our model that the principal and platform know workers’ outcome distributions, we note that crowdsourcing platforms systematically track task completion rates and historical outcomes. This enables them to estimate outcome distributions based on accumulated data. While the principal lacks direct access to this data, they can still infer outcome distributions from publicly available information: Platforms display ratings, completion rates, and certifications, and workers often showcase their past successes in their public profiles to attract employers. These signals provide the principal with a reasonable estimate of success likelihood. It is worth emphasizing that all our results also hold when the outcome distributions are private (see Appendix A for a detailed discussion of this assumption).
@@ 2.3 Revenue-Maximization Mechanism Design
In this subsection, we revisit the single-item revenue maximization in the Bayesian environment. First, we consider the scenario where a seller intends to sell a single item to a single buyer. The buyer’s private value of the item, denoted by $v$, is drawn from a prior distribution $F$ [^10] With a slight abuse of notation, we also use $F$ and $F^{\prime}$ to denote the cumulative distribution function (CDF) and probability density function (PDF) of the distribution $F$, respectively. In addition, we denote by $\mathtt{supp}(F)$ as the support of distribution $F$.
@@
Additionally, the cumulative hazard rate function $H\colon\mathtt{supp}(F)\to\mathbb{R}_{\geq 0}$ is defined as $H(p)\triangleq\int_{0}^{p}h(p)=-\ln(1-F(p))$. In the following, we introduce two important types of distributions.
Regular distributions.  A distribution $F$ is regular if the corresponding virtual value function $\phi(\cdot)$ is monotone non-decreasing in $\mathtt{supp}(F)$. Assuming that the distribution $F$ is regular, $\phi^{-1}(0)$ [^12] Note that the virtual value function $\phi(\cdot)$ may not be strictly increasing, i.e., there can be two different values both corresponds to the same virtual value. For ease of notation, throughout the paper, we always break ties by choosing the largest one, namely, $\phi^{-1}(p)=\max\left\{z\geq 0\,:\,\phi(z)\leq p\right\}$. is exactly the monopoly reserve price $\theta$. Furthermore, due to the monotonicity of $\phi(\cdot)$ and the definition of the monopoly reserve price $\theta$, we have $\phi(p)\geq 0$ for all $p\geq\theta$.
@@
Sometimes, it is more convenient to work in the quantile space rather than in the actual value domain. Specifically, the quantile of a distribution $F$ with respect to a price $p\in\mathtt{supp}(F)$ is defined as $q\triangleq 1-F(p)$. Using this, the revenue curve $R\colon[0,1]\to\mathbb{R}_{\geq 0}$ of the distribution $F$ is defined as $R(q)\triangleq q\cdot F^{-1}(1-q)$ for any $q\in[0,1]$. In other words, if $p\in\mathtt{supp}(F)$ is a price and $q=1-F(p)$ is the corresponding quantile, $R(q)$ represents the expected revenue of selling the item at price $p$. It can be verified that the derivative $R^{\prime}(q)$ equals the virtual value $\phi(p)$, where $p=F^{-1}(1-q)$. The regularity of $F$ implies that $R^{\prime}(\cdot)$ is non-increasing, making $R(\cdot)$ concave. Additionally, the MHR property of $F$ implies that the cumulative hazard rate function $H(\cdot)$ is convex.
Now, we proceed to consider the multi-buyer scenario where a seller wants to sell a single item to one of $n$ buyers, whose private values of the item, $\left\{v_{i}\right\}_{i\in[n]}$, are drawn independently from regular distributions $\left\{F_{i}\right\}_{i\in[n]}$. Myerson (1981) proved that the virtual welfare maximizer mechanism is DSIC, ex-post IR and maximizes the seller’s expected revenue.
Definition 2.3 (Virtual welfare maximizer, Myerson, 1981).
  1. 1.
  2. 2.
  3. 3.
    $p_{i}^{{\dagger}}(\bm{v})\triangleq\phi_{i}^{-1}\left(\max\nolimits_{j\in[n]\setminus\left\{i\right\}}{\left(\phi_{j}\left(v_{j}\right)\right)^{+}}\right)\cdot x^{{\dagger}}_{i}(\bm{v})$
Lemma 2.1 (Myerson, 1981).
@@ 2.4 Assumptions and Notations
The first three properties are standard in the mechanism design literature. In particular, by applying the standard ironing technique (Myerson, 1981), the main results in Sections 3 and 4 can be extended to scenarios where the cost distribution is irregular. The final property holds without loss of generality, as an agent will never be selected to complete the task if her realized contribution is always non-positive.
We use the notation $[n]=\{1,2,\ldots,n\}$ and $[s:\ell]=\{s,s+1,\ldots,\ell\}$ for $s\leq\ell$. Given a distribution $F$, we reuse $F$ to denote its CDF and let $F^{\prime}$ represent its PDF. Bold symbols are used to denote vectors or joint distributions. For instance, $\bm{v}=(v_{1},v_{2},\ldots,v_{n})$ and $\bm{F}=F_{1}\otimes F_{2}\otimes\cdots\otimes F_{n}$. Specifically, for $n$ identical distributions $\left\{F\right\}_{i=1}^{n}$, we denote the joint distribution $\bm{F}$ as $F^{\otimes n}$. For $n$ random variables $\bm{v}\sim\bm{F}$, we use $v_{\max}$ to denote the first-order statistic among them, i.e., $v_{\max}\triangleq\max_{i\in[n]}v_{i}$. For any $z\in\mathbb{R}$, define ${\left(z\right)^{+}}\triangleq\max\left\{z,0\right\}$.
3 Optimal Mechanism for the Intermediary
In this section, we assume that a contract has been issued by the principal to the intermediary and focus on the mechanism design problem between the intermediary and $n$ agents.
Our main result, Theorem 3.1, demonstrates that the optimal mechanism for the intermediary is essentially the virtual welfare maximizer mechanism, as described in Definition 2.3, where the intermediary and the agents correspond to the seller and the buyers, respectively, and the agents’ contracted contributions are equivalent to the buyers’ values of the auctioned item.
Theorem 3.1 (Optimal mechanism characterization).
    $\bm{w}_{\bm{t}}=\left(\left\langle\bm{t},\bm{\gamma}_{1}\right\rangle-c_{1},\left\langle\bm{t},\bm{\gamma}_{2}\right\rangle-c_{2},\ldots,\left\langle\bm{t},\bm{\gamma}_{n}\right\rangle-c_{n}\right)$
  • If above virtual values are all negative, no agent is assigned the task; otherwise, assign the task to the agent with the maximum virtual value.
    $p_{\bm{t},i}^{\star}(\bm{c})\triangleq\left(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-\phi_{\bm{t},i}^{-1}\left(\max_{j\in[n]\setminus\left\{i\right\}}{\left(\phi_{\bm{t},j}\left(w_{\bm{t},j}\right)\right)^{+}}\right)\right)\cdot x^{\star}_{\bm{t},i}(\bm{c})$
An immediate implication from above result is as follows.
Corollary 3.2.
The proof of Theorem 3.1 is relatively standard and thus deferred to Section C.1. At a high level, we first show that any truthful mechanism between the intermediary and $n$ agents can be converted into a truthful single-item auction between a seller and $n$ buyers, and vice versa, as described in Lemma C.1. The key is to convert agent $i$’s private cost $c_{i}$ into buyer $i$’s private value $v_{i}$ in the auction. The contract $\bm{t}$ plays a pivotal role in this conversion: $v_{i}=\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i}$, which is the contracted contribution of agent $i$. Thus, the seller’s revenue in the auction meets the intermediary’s utility in the mechanism. Then, we establish that the regularity or MHR property of the contribution distribution is preserved in the distribution of the contracted contribution for any contract $\bm{t}$, as shown in Lemma C.2. This crucial property enables us to apply standard auction theory, demonstrating that the optimal mechanism for the intermediary is indeed the virtual welfare maximizer mechanism, as detailed in Definition 2.3 and Lemma 2.1.
@@ 4 Optimal Contract for the Principal
In this section, we use backward induction to characterize the optimal contract for the principal.
The main theorem of this section characterizes the principal’s problem of designing an optimal contract to maximize her utility in various scenarios.
@@
    For general instances, the optimal contract and the corresponding utility of the principal is the optimal solution and objective value of the optimization defined as follows:
@@
    For instances with identical expected reward profiles, i.e., $\bm{\gamma}_{i}\equiv\bm{\gamma}$ for each $i\in[n]$, there exists an optimal contract that is linear. In particular, this optimal linear contract and the corresponding utility of the principal is the optimal solution and objective value of the optimization defined as follows:
@@
    For instances with identical expected reward profiles and contribution distributions [^13] This is equivalent to assuming identical outcome distributions and cost distributions., i.e., $\bm{\gamma}_{i}\equiv\bm{\gamma}$ and $F_{i}\equiv F$ for each $i\in[n]$, there exists an optimal contract that is linear. In particular, the optimal utility of the principal is the optimal objective value of the optimization program defined as follows:
@@
We first discuss the implication of the above characterization and then sketch its proof. The formal proof can be found in Section C.2.
@@
Implications of Theorem 4.1 The first two formulations in Theorem 4.1 address scenarios with heterogeneous agents (i.e., agents differ in either rewards or cost distributions), and are expressed as summations over all agents, with each term representing the utility an agent brings to the principal multiplied by that agent’s probability of winning. In contrast, the third formulation considers essentially homogeneous agents (i.e., all agents share identical reward and cost distributions), and is expressed as the utility a single agent provides to the principal multiplied by the probability that at least one agent is selected. Next, we discuss each of these formulations separately.
@@
The first formulation reformulates the principal’s problem by expressing each agent’s winning probability as a function of the contract $\bm{t}$. Recall each agent’s contracted contribution is determined by $w_{\bm{t},i}=\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i}$, which depends on the agent’s cost, expected reward profile, and the chosen contract. Consequently, different contracts for distinct outcomes can reshape the ranking of agents’ contracted contributions, thereby affecting the selection of winners. Furthermore, since winning probabilities are tied to the distributions of contracted contributions, they become contract-dependent as well.
@@
The second formulation assumes all agents have the same expected reward profiles $\bm{\gamma}_{i}\equiv\bm{\gamma}$ but can differ in cost distributions. Under this condition, we show the optimality of linear contract, thereby simplifying the principal’s contract design space from the $m$-dimensional space (i.e., $\mathbb{R}_{+}^{m}$) to the single-dimensional space (i.e., $[0,1]$). Notably, the principal’s contract design problem can be simplified to setting a “discriminatory threshold” $z_{\alpha,i}\triangleq(1-\alpha)r+\theta_{\alpha,i}$ for each agent $i$. Agent $i$ wins if her contribution is the highest and meets or exceeds her posted price. The principal’s utility from agent $i$ is determined by the virtual value of her threshold $\phi_{i}(z_{\alpha,i})$ (instead of the virtual value of her contribution). Since $z_{\alpha,i}=(1-\alpha)r+\theta_{\alpha,i}$ is defined for each agent $i\in[n]$, the discriminatory threshold profile $\left\{z_{\alpha,i}\right\}_{i=1}^{n}$ is restricted by $\alpha$. Changes to the contract do not affect the ranking of the agents’ private values or the identity of the winner; the winner remains fixed across all contracts, given realizations of agents’ costs. While the threshold for each agent depends on the contract, her winning probability remains unchanged as long as her contribution meets or exceeds her posted threshold.
@@
The third formulation assumes all agents have identical expected reward profiles and contribution distributions, simplifying the principal’s problem further into a virtual value pricing problem. Here, the principal posts an anonymous price $z$. A sale occurs if some agent’s contribution meets or exceeds this price. The principal realizes a utility equal to the virtual value of the posted price $z$ instead of the price itself.
@@
Proof overview of Theorem 4.1 We outline the main ideas behind the proof of Theorem 4.1. To obtain the first formulation for the general case, we begin by utilizing the results from Theorem 3.1 to express each agent’s winning probability. This allows us to reformulate the principal’s maximization problem as maximizing the sum of each agent’s winning probability times the principal’s utility brought by her. For the case of identical expected reward profiles, we analyze the relationship between each agent’s contribution and their value. Then we convert both the winning probabilities and the principal’s utility into expressions related to the contributions and their corresponding virtual values. This conversion relies on the condition that expected reward profiles are identical. To establish the third formulation for the identical expected reward profiles and distributions case, we reformulate the sum of winning probabilities to represent the probability that the first-order statistic of the agents’ contribution meets or exceeds the anonymous price set by the contract. Finally, we establish that the principal’s maximum utility from designing an optimal contract aligns with the maximum revenue achievable by selecting an optimal price from the support of the contribution distribution, as demonstrated by analyzing the value range of the anonymous price derived from the contract (see Lemma 4.2 and its proof in Section C.3).
Lemma 4.2 (Continuity of monopoly reserve).
For any fixed agent $i\in[n]$, consider the monopoly reserve price $\theta_{\alpha,i}$ as a function of the linear contract $\alpha$. Then, $\theta_{\alpha,i}$ is continuous in w.r.t. $\alpha\in[0,1]$.

### v2

@@ 2.2 Model Discussion
In this subsection, we discuss key assumptions underlying modeling choices.
To justify the assumption that the principal initiates the contract, we note that in real-world crowdsourcing markets, requesters typically hold greater bargaining power because multiple intermediaries compete for their business. For example, both Upwork and Freelancer charge clients a similar fee of around 4% (Mert Bulut, 2024). This competitive dynamic is further illustrated by shifts in the microtask market: on Amazon Mechanical Turk, requesters historically benefited from low fees, but after Amazon raised its commission to 20% in 2015, many migrated to alternatives such as Prolific, which offered better user terms and higher data quality (Banakou and Slater, 2023). These observations align with standard contract theory, where principals with strong outside options propose contracts that agents accept.
@@
To justify the assumption in our model that the principal and platform know workers’ outcome distributions, we note that crowdsourcing platforms systematically track task completion rates and historical outcomes. This enables them to estimate outcome distributions based on accumulated data. While the principal lacks direct access to this data, she can still infer outcome distributions from publicly available information: Platforms display ratings, completion rates, and certifications, and workers often showcase their past successes in their public profiles to attract employers. These signals provide the principal with a reasonable estimate of success likelihood. It is worth emphasizing that all our results continue to hold when the outcome distributions are private (see Appendix A for a detailed discussion of this assumption).
@@ 2.3 Revenue-Maximization Mechanism Design
In this subsection, we revisit the single-item revenue maximization in the Bayesian environment. We consider the scenario where a seller intends to sell a single item to a single buyer. The buyer’s private value of the item, denoted by $v$, is drawn from a prior distribution $F$.  [^11] With a slight abuse of notation, we also use $F$ and $F^{\prime}$ to denote the cumulative distribution function (CDF) and probability density function (PDF) of the distribution $F$, respectively. Besides, $\mathtt{supp}(F)$ denotes the support of $F$.
@@
Additionally, the cumulative hazard rate function $H\colon\mathtt{supp}(F)\to\mathbb{R}_{\geq 0}$ is defined as $H(p)\triangleq{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\int_{0}^{p}h(x){\mathrm{d}}x}=-\ln(1-F(p))$. In the following, we introduce two important types of distributions.
Regular distributions. A distribution $F$ is regular if the corresponding virtual value function $\phi(\cdot)$ is monotone non-decreasing in $\mathtt{supp}(F)$. Assuming that the distribution $F$ is regular, $\phi^{-1}(0)$ [^13] Note that the virtual value function $\phi(\cdot)$ may not be strictly increasing, i.e., there can be two different values both corresponding to the same virtual value. For ease of notation, throughout the paper, we always break ties by choosing the largest one, namely, $\phi^{-1}(p)=\max\left\{z\geq 0\,:\,\phi(z)\leq p\right\}$. is exactly the monopoly reserve price $\theta$. Furthermore, due to the monotonicity of $\phi(\cdot)$ and the definition of the monopoly reserve price $\theta$, we have $\phi(p)\geq 0$ for all $p\geq\theta$.
@@
Quantile space, revenue curve, and cumulative hazard rate function. Sometimes, it is more convenient to work in the quantile space rather than in the actual value domain. Specifically, the quantile of a distribution $F$ with respect to a price $p\in\mathtt{supp}(F)$ is defined as $q\triangleq 1-F(p)$. Using this, the revenue curve $R\colon[0,1]\to\mathbb{R}_{\geq 0}$ of the distribution $F$ is defined as $R(q)\triangleq q\cdot F^{-1}(1-q)$ for any $q\in[0,1]$. In other words, if $p\in\mathtt{supp}(F)$ is a price and q=1F(p)q=1-F(p) is the corresponding quantile, $R(q)$ represents the expected revenue of selling the item at price $p$. It can be verified that the derivative $R^{\prime}(q)$ equals the virtual value $\phi(p)$, where $p=F^{-1}(1-q)$. The regularity of $F$ implies that $R^{\prime}(\cdot)$ is non-increasing, making $R(\cdot)$ concave. Additionally, the MHR property of $F$ implies that the cumulative hazard rate function $H(\cdot)$ is convex.
@@ 2.4 Assumptions and Notations
The first three properties are standard in the mechanism design literature. In particular, by applying the standard ironing technique (Myerson, 1981), the main results in Section 3 can be extended to scenarios where the cost distribution is irregular. The final property holds without loss of generality, as an agent will never be selected to complete the task if her realized contribution is always non-positive.
We use the notation $n]=\{1,2,\ldots,n\}$ and $s:\ell]=\{s,s+1,\ldots,\ell\}$ for $s\leq\ell$. Given a distribution $F$, we reuse $F$ to denote its CDF and let $F^{\prime}$ represent its PDF. Bold symbols are used to denote vectors or joint distributions. For instance, $\bm{v}=(v_{1},v_{2},\ldots,v_{n})$ and $\bm{F}=F_{1}\otimes F_{2}\otimes\cdots\otimes F_{n}$. Specifically, for $n$ identical distributions $\left\{F\right\}_{i=1}^{n}$, we denote the joint distribution $\bm{F}$ as $F^{\otimes n}$. For $n$ random variables $\bm{v}\sim\bm{F}$, we use $v_{\max}$ to denote the first-order statistic among them, i.e., $v_{\max}\triangleq\max_{i\in[n]}v_{i}$. For any $z\in\mathbb{R}$, define ${\left(z\right)^{+}}\triangleq\max\left\{z,0\right\}$.
@@ 3 Optimal Contract for the Principal
In this section, we characterize the optimal contract for the principal under various scenarios.
@@
    For general instances, the optimal contract and the corresponding payoff of the principal is the optimal solution and objective value of the optimization defined as follows:
@@
    For instances with identical expected reward profiles (i.e., $\bm{\gamma}_{i}\equiv\bm{\gamma}$ with each $i\in[n]$), there exists an optimal contract that is linear. In particular, this optimal linear contract and the corresponding payoff of the principal are the optimal solution and objective value of the optimization defined as follows:
@@
    For instances with identical expected reward profiles and contribution distributions [^14] This is equivalent to assuming identical outcome distributions and cost distributions. (i.e., $\bm{\gamma}_{i}\equiv\bm{\gamma}$ and $F_{i}\equiv F$ for each $i\in[n]$), there exists an optimal contract that is linear. In particular, the optimal payoff of the principal is the optimal objective value of the optimization program defined as follows:
@@
We first discuss the implications of the above characterization and then sketch its proof. The formal proof can be found in Section C.2.
@@
Implications of Theorem 3.1. The first two formulations in Theorem 3.1 address scenarios with heterogeneous agents (i.e., agents differ in either rewards or cost distributions), and are expressed as summations over all agents, with each term representing the utility an agent brings to the principal multiplied by that agent’s probability of winning. In contrast, the third formulation considers essentially homogeneous agents (i.e., all agents share identical reward and cost distributions), and is expressed as the utility a single agent provides to the principal multiplied by the probability that at least one agent is selected. Next, we discuss each of these formulations separately.
@@
The first formulation reformulates the principal’s problem by expressing each agent’s winning probability as a function of the contract $\bm{t}$. Recall that each agent’s contracted contribution is determined by $w_{\bm{t},i}=\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-c_{i}$, which depends on the agent’s cost, expected reward profile, and the chosen contract. Consequently, different contracts for distinct outcomes can reshape the ranking of agents’ contracted contributions, thereby affecting the selection of winners. Furthermore, since winning probabilities are tied to the distributions of contracted contributions, they become contract-dependent as well.
@@
The second formulation assumes all agents have the same expected reward profiles $\bm{\gamma}_{i}\equiv\bm{\gamma}$ but can differ in cost distributions. Under this condition, we show the optimality of the linear contract, thereby simplifying the principal’s contract design space from the $m$-dimensional space (i.e., $\mathbb{R}_{+}^{m}$) to the single-dimensional space (i.e., $[0,1]$). Notably, the principal’s contract design problem can be simplified to setting a “discriminatory threshold” $z_{\alpha,i}\triangleq(1-\alpha)r+\theta_{\alpha,i}$ for each agent $i$. Agent $i$ wins if her contribution is the highest and meets or exceeds her posted price. The principal’s utility from agent $i$ is determined by the virtual value of her threshold $\phi_{i}(z_{\alpha,i})$ (instead of the virtual value of her contribution). Since $z_{\alpha,i}=(1-\alpha)r+\theta_{\alpha,i}$ is defined for each agent $i\in[n]$, the discriminatory threshold profile $\left\{z_{\alpha,i}\right\}_{i=1}^{n}$ is restricted by $\alpha$. Consequently, conditional on a realized cost profile, changing the contract $\alpha$ does not affect the ranking of the agents’ private values or the identity of the winner. While the threshold for each agent depends on the contract, her winning probability remains unchanged as long as her contribution meets or exceeds her posted threshold.
@@
The third formulation assumes all agents have identical expected reward profiles and contribution distributions, simplifying the principal’s problem further into a virtual value pricing problem. Here, the principal posts an anonymous price $z$. A sale occurs if at least one agent’s contribution meets or exceeds this price. The principal realizes a utility equal to the virtual value of the posted price $z$ instead of the price itself.
@@
Proof overview of Theorem 3.1. We outline the main ideas behind the proof of Theorem 3.1. To derive the first formulation in the general cases, we begin by applying the characterization of the intermediary’s optimal mechanism from Proposition 3.2 to express each agent’s winning probability. This enables us to reformulate the principal’s problem as maximizing, over all contracts, the sum over agents of their winning probabilities weighted by the principal’s utility from each agent. For the case in which agents have identical expected reward profiles, we analyze the relationship between each agent’s contribution and her value. Then we rewrite both the winning probabilities and the principal’s utility in terms of agents’ contributions and the corresponding virtual values. This conversion relies on the assumption that expected reward profiles are identical. To establish the third formulation, which additionally assumes identical contribution distributions, we reformulate the sum of winning probabilities as the probability that the first-order statistic of the agents’ contribution exceeds the anonymous price induced by the contract. Finally, we show that the principal’s maximum expected utility from designing an optimal contract aligns with the maximum revenue obtainable by selecting an optimal price from the support of the contribution distribution, as demonstrated by analyzing the value range of the anonymous price derived from the contract (see Lemma C.5).
Proposition 3.2 (Optimal mechanism characterization).
    $\bm{w}_{\bm{t}}=\left(\left\langle\bm{t},\bm{\gamma}_{1}\right\rangle-c_{1},\left\langle\bm{t},\bm{\gamma}_{2}\right\rangle-c_{2},\ldots,\left\langle\bm{t},\bm{\gamma}_{n}\right\rangle-c_{n}\right),$
  • If the above virtual values are all negative, no agent is assigned the task; otherwise, assign the task to the agent with the maximum virtual value.
    $p_{\bm{t},i}^{\star}(\bm{c})\triangleq\left(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle-\phi_{\bm{t},i}^{-1}\left(\max_{j\in[n]\setminus\left\{i\right\}}{\left(\phi_{\bm{t},j}\left(w_{\bm{t},j}\right)\right)^{+}}\right)\right)\cdot x^{\star}_{\bm{t},i}(\bm{c}),$
Proposition 3.2 demonstrates that the optimal mechanism for the intermediary is essentially the virtual welfare maximizer mechanism, as described in Definition C.1, where the intermediary and the agents correspond to the seller and the buyers, respectively, and the agents’ contracted contributions are equivalent to the buyers’ values of the auctioned item. The proof of Proposition 3.2 is relatively standard and thus deferred to Section C.1.